Angles A and B are complementary angles. If the measure o angle A is 30 more than the measure of angle B. Find the measures of the two angles
Answer by IWork4Dessert(60)

Complementary angles are angles whose values, when added together, equal 90º.
Let angle A equal "x+30" and angle B equal "x", as the instructions state that angle A is thirty more than angle B.
Now you create an equation using this.
(x)+(x+30)=90
Combine your like terms, so your variables(x).
2x+30=90
Subtract the 30 from both sides, since the 30 is positive, to cancel it out.
2x=60
Divide both sides by 2 since the x has a coefficient, or number in front of it. You want to get the variable totally alone on one side of the equation.
x=30
Now check the answer by putting it back into the equation.
(30)+(30+30)=90
30+60=90
90=90
There's your answer.
Hope this helps!